OpenStudy (kayluh123):

There are 45 new houses being built in a neighborhood. Last month, 1/3 of them were sold. This month, 1/5 of the remaining houses were sold. How many houses are left to be sold?

OpenStudy (calculusxy):

1/3 x 45 = 15 45-15=30 1/5 x 30 = 6 30-6=24 So 24 houses are left to be sold.

OpenStudy (anonymous):

1) 1/3 of 45 is 15 (this can be found by simply dividing 45 by 3. 2) Subtract this amount from the total, so that 45-15 = 30. 3) 1/5 of the remaining houses being sold would be 30 / 5, or 6. 4) Once again, subtract this amount from the total, resulting in 24 houses left.